Determine whether each of these functions is a bijection from R to R (real).a) f (x) = 2x + 1b) f (x) = x2 + 1c) f (x) = x3

1
Expert's answer
2022-01-06T18:01:58-0500

Let us determine whether each of these functions is a bijection from "\R" to "\R."


a) "f (x) = 2x + 1"


Let "f(x_1)=f(x_2)." Then "2x_1+1=2x_2+1," and thus "2x_1=2x_2." We conclude that "x_1=x_2," and hence the function "f" is one-to-one.

For any "y\in\R" there exists "x=\frac{y-1}2\in\R" such that "f(x)=f(\frac{y-1}2)=2\frac{y-1}2+1=y-1+1=y,"

and hence the function "f" is a surjection.

Consequently, the function "f" is a bijection.


b) "f (x) = x^2 + 1"


Since "f (-1) = (-1)^2 + 1=1^2+1=f(1)," we conclude that this function is not one-to-one, and therefore "f" is not a bijection.


c) "f (x) = x^3"


Let "f(x_1)=f(x_2)." Then "x_1^3=x_2^3," and thus "x_1=x_2." We conclude that the function "f" is one-to-one.

For any "y\in\R" there exists "x=\sqrt[3]{y}\in\R" such that "f(x)=f(\sqrt[3]{y})=(\sqrt[3]{y})^3=y,"

and hence the function "f" is a surjection.

Consequently, the function "f" is a bijection.